# Jie Wu1∗  Haoling Li1∗  Xin Zhang2∗†‡  Jiani Guo3  Jane Luo2  Steven Liu2 Yangyu Huang2Ruihang Chu1‡Scarlett Li2Yujiu Yang1

### v1

@@ 3 Experiment
Setup. In this study, we adopt GPT-o3-mini (OpenAI, 2025) for task formulation, Deepseek-R1-0528 (DeepSeek-AI, 2025) and Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025) for solution sampling, and R1-0528 for test case generation. Statistics for SFT datasets are provided in §C.2. For SFT, we set the learning rate at 5e-5, with a global batch size of 128 to train 8 epochs. For RL, the reward is defined as the fraction of passed tests among all given tests (detailed in §A.2). The program executes in an isolated sandbox environment deployed with Redis, which supports optimized concurrent code testing (infrastructure details are provided in §A.5). Training configurations and costs are supplemented in §A.4.
Evaluation. We evaluate Code LLMs on LiveCodeBench (Jain et al., 2024) v5 (covering problems released between Aug. 2024 and Feb. 2025) and v6 (Feb. to May 2025), which are the most widely used benchmarks for code reasoning models. Baselines are documented in §A.6. To ensure a fair comparison, we use Qwen2.5-Coder-7B-Instruct and Qwen3-8B-Base as backbones, and report the avg@8 pass rate using a sampling temperature of 0.6 with top-p 0.95 to align with the baselines.
@@
MiMo-SFT (Xiaomi et al., 2025) MiMo-Base 7B 500k Unclear avg@8 52.3 45.5
@@
[…]
SFT-then-RL Baselines (Stage 2)
AceReason1.1 AceReaon1.1-SFT 7B - Real avg@8 57.2 52.1
MiMo MiMo-SFT 7B 130k Unclear avg@8 57.8 49.3
Klear-Reasoner Klear-Reasoner-SFT 8B 106k Real avg@8 61.6 53.1
[…]
3.1 Main Results
As shown in Table 1, during the SFT stage, X-Coder-SFT achieves an avg@8 pass rate of 60.3. Compared with RL baselines, X-Coder-SFT exhibits a clear advantage over 14B-based RL models (e.g., DeepCoder-Preview-14B, AReal-boba²-14B), despite those models being built on the stronger foundation R1-Distilled-Qwen. Relative to SFT-then-RL models, X-Coder further boosts its performance after RL, reaching 62.9. On Qwen3-Base, X-Coder attains an avg@8 pass rate of 64.0.
3.2 SFT Experiments and Analysis
During the SFT stage, we investigate a central question: Can the SFT dataset be effectively scaled, and along which dimension should it be scaled more favorably? To explore this, we are inspired by AceReason-Nemotron 1.1 (Liu et al., 2025b) and expand the SFT dataset from two distinct perspectives: increasing the number of unique tasks and enlarging the number of solutions per task. We design seven subsets (v1–v6): v1–v4 increase the number of unique tasks (32k, 64k, 128k, and 192k unique prompts, each with 1 solution), while v5–v6 expand the number of solutions per task (16k unique prompts with 4 solutions, and 8k unique prompts with 8 solutions). The results in Figure 3 reveal a promising scaling trend, where v4 >> v3 >> v2 >> v1, with performance steadily improving from 43.7% to 62.7%.
Figure 3: Scaling laws on the SFT dataset generated by SynthSmith. Left: Performance comparison of on LiveCodebench v5 to examine scaling trend. Right: Performance comparison across scaling unique tasks and scaling solutions per task.
[…]
Comparison with Real-World and Synthetic Code Datasets. For real-world datasets, we compare against OpenCodeReasoning (Ahmad et al., 2025), the largest reasoning-based synthetic dataset to date for competitive coding. We train our dataset and OpenCodeReasoning using the same number of training tokens with Qwen2.5-Coder-7B-Instruct. The results are shown in Table 2. Our proposed dataset yields a 6.7-point improvement after SFT, with most gains coming from the medium and hard splits. The improvement is attributed to our pipeline’s ability to synthesize more challenging tasks, which demand longer reasoning (average length 17.7k vs. 8.0k), and to provide greater prompt diversity, which proves more effective than increasing solution diversity.
Table 2: OpenCodeReasoning vs. Dataset from SynthSmith.
Model Avg. Easy Medium Hard
OCR-Qwen-7B-Instruct (Ahmad et al., 2025) 51.3 95.4 64.0 18.0
OCR-Qwen-Coder-7B-Instruct 53.6 95.2 67.0 21.8
X-Coder-Qwen-Coder-7B-Instruct 60.3 (+6.7) 96.8 73.3 37.8
Table 3: Synthetic Data by SelfCodeAlign vs. by SynthSmith.
Method Task Gen. Ans. Gen. Data Score
SelfCodeAlign (Wei et al., 2024) GPT-o3-mini DeepSeek-R1 10k 27.1
SynthSmith (Ours) GPT-o3-mini DeepSeek-R1 10k 31.7 (+4.6)
For synthetic datasets, we implemented the SelfCodeAlign (Wei et al., 2024) method using same teacher models and adapted it to the competitive programming domain to deliver a 10k-sample dataset. The results in Table 3 shows our method achieves a 4.6 performance gains, demonstrating the effectiveness of our data synthesis strategy for competitive programming.
@@ 3.3 RL Experiments and Analysis
Our investigation of the RL stage uncovers the following key insights into its role and behavior:
(i) RL as a Powerful Refiner. RL fine-tuning is not merely an incremental add-on but a powerful optimization step. As shown in Table 1, when applied to a converged SFT model using only code data, it yields a substantial 4.6% absolute gain in average pass-rate. This highlights RL’s unique capability to refine policy beyond the distribution of the initial supervised dataset.
(ii) The “Good-gets-Better” Principle. RL performance is tightly coupled to the strength of the SFT initializer. Using two SFT models trained on similar data distributions but with different LiveCodeBench scores as starting points, we observe in Figure 4 that, under identical RL settings, the stronger initializer consistently attains higher rewards.
Figure 4: Reward comparison of weak and strong SFT models as RL initializer.
A stronger SFT foundation enablings to explore a more promising policy space and achieve a higher performance ceiling. This underscores the importance of a high-quality initial model as a prerequisite for effective RL.
(iii) Resilience to Noisy Supervision. Contrary to the common assumption that RL requires pristine reward signals, our experiments reveal a resilience to data imperfections during RL. The model also effectively benefits from synthetic test cases, suggesting that RL can be successfully deployed in scenarios with large-scale but imperfect feedback (Wang et al., 2020; Lv et al., 2025), significantly lowering the barrier to code RL data collection.
@@ 4 Ablation Study
Despite the strong performance of X-Coder, the determinants of high-quality synthetic data for SFT remain insufficiently understood. To elucidate these factors, we conduct a comprehensive ablation along six axes: (i) the effect of the proposed dual-verification strategy; (ii) the impact of distinct thinking types in solutions; (iii) the influence of task styles; (iv) a head-to-head comparison of tasks produced by SynthSmith versus those from open-source synthetic datasets; (v) data-selection strategies to identify patterns that shape downstream performance; and (vi) comparison of prompting-based and tool-based test generation strategies.
(a) Raw vs. verified solutions.
(b) Task style comparison.
(c) Comparison with EpiCoder.
Figure 5: Ablations on verification, task style, and task sources.
Q1: Dual-verification for High-Quality Data Curation. To mitigate the noise introduced by stochastically sampled solutions, we employ a dual-verification strategy for data curation. This strategy first leverages the self-consistency principle to identify the most likely correct solution from multiple candidates. Subsequently, these candidate solutions are executed against a comprehensive set of test cases to verify their functional correctness and robustness, thereby capturing subtle runtime errors (e.g., ValueError, IndexError, or Timeout) that are undetected by static analysis methods like AST checks. The efficacy of this approach is validated by our empirical results, as shown in Figure 5(a). Using an identical backbone (Qwen2.5-Coder-7B-Instruct) and dataset (64k tasks), the model trained on verified solutions significantly outperforms its counterpart trained on raw solutions. However, this quality assurance comes at a considerable computational cost. For instance, fully verifying 200k samples necessitates the generation of 1.6 million long-CoT trajectories and 24 million test executions. This overhead establishes a clear trade-off, as prior work (Li et al., 2025; Gandhi et al., 2025) indicates that models can still learn effectively from unverified long-CoT data, making raw-solution training a more resource-efficient, albeit potentially less performant, alternative.
Q2: Solution Types: Long CoT vs. Short CoT. The length of CoT proves to be a critical factor for performance, with longer CoTs yielding superior results despite higher training costs. To demonstrate this, we compare the Qwen2.5-Coder-7B-Instruct trained on solutions generated by DeepSeek-R1-0528 (Long-CoT) and Qwen3-235B-A22B-Instruct-2507 (Short-CoT) for an identical set of tasks (200k).
@@
Q3: Ablation on Task Style. We evaluate the effect of task styles (AtCoder, Codeforces, and LeetCode) by synthesizing three corpora of 32k tasks each (8k unique problems with 4 solutions per problem) from identical input features. For each corpus, solutions are generated with DeepSeek-R1-0528 and used to fine-tune the Qwen2.5-Coder-7B-Instruct. Results are shown in Figure 5(b). Although AtCoder-style tasks yield slightly higher scores, we adopt Codeforces-style as the predominant format in our demonstration dataset (Codeforces : AtCoder : LeetCode = 70 : 15 : 15), reflecting its prominence as the mainstream competitive-programming platform.
Q4: Tasks from SynthSmith vs Tasks from EpiCoder-380k. We randomly select 64k tasks from our SFT dataset and another 64k from EpiCoder-380k, and use DeepSeek-R1-0528 to complete solutions. Figure 5(c) shows that tasks from SynthSmith yield a 21% absolute performance gain, demonstrating its ability to produce high-quality tasks tailored for competitive programming.
@@
Q5: Data Selection. To investigate data utilization efficiency, we explore task selection strategies for competitive programming. Specifically, we evaluate three approaches: (1) difficulty-based selection, where GPT-4o-2411 assigns discrete difficulty scores to tasks, simulating the Codeforces rating system; (2) rationale-based selection, where DeepSeek-R1-0528 generates CoT reasoning for each task, and tasks that elicit longer reasoning traces are prioritized; and (3) random selection as a baseline. For validation, each strategy independently samples a 50k-task subset from a 200k-task pool. Solutions are generated by Qwen3-235B-A22B-Instruct-2507, and models were trained for three epochs with a 16k context length.
As shown in Figure 6, tasks that induce longer CoT are regarded as more valuable training data for competitive programming, as they demand deeper reasoning and are potentially more challenging.
Q6: Prompting-based vs. Tool-based Test Generation. We compare prompting-based and tool-based test generation using tasks from CodeContests (Li et al., 2022). We leverage the corresponding golden solutions to evaluate the accuracy and complexity of the tests produced by the two approaches. The results in Table 5 show that the tool-based approach outperforms the prompting-based method across multiple dimensions. Qualitatively, it is more versatile, capable of systematically generating random, scalable, boundary, and stress tests, which are essential for robust code evaluation but not supported by prompting-based methods.
Prompting-based low 13.6 5 15 82.0% 77.4%
Tool-based high 18.3 5 27 78.8% 87.9%
5 Discussion
In this section, we present an in-depth analysis of the main challenges in code reasoning. Reasoning models often suffer from assertion errors, highlighting persistent reasoning limitations on harder tasks. We further identify a mediation pattern among task difficulty, reasoning length, and pass rate, and extend our analysis with test-time scaling experiments and case studies on cognitive behavior, reward hacking, and undesirable patterns.
Failure Analysis. We classify failure cases into seven types: Wrong Answer (output mismatches the expected answer), Time Limit Exceeded, Memory Limit Exceeded, No Code Block Generated (truncated due to heavy reasoning before the final code is generated), Incomplete Code Block (partial code without closure), Function Signature Mismatch (incorrect function definition), and Syntax Error (complete code with syntax issues). The error distribution in Table 6 indicates that the primary bottleneck lies in reasoning capability, with most errors stemming from wrong answers. Two other major failure categories are No Code Block Generated and Time Limit Exceeded (TLE). We carefully inspected the no-code samples and found that all of them exceeded the 32k context window, causing the reasoning process to be truncated and incomplete. The frequency of TLE errors highlights the need for Code LLMs to prioritize code efficiency.
Table 6: Distribution of failure cases for 16 rollouts on LiveCodeBench v5 (268 tasks).
@@ 5 Discussion
Pass Rate by Reasoning Token Length. The results in Table 7 show that the pass rate decreases sharply as reasoning token length increases, exhibiting a clear downward trend. This finding runs counter to the intuitive expectation that greater test-time token usage reflects deeper reasoning and should therefore yield higher accuracy. Instead, we observe a significant chained relationship among problem difficulty, reasoning length, and pass rate: problem difficulty is positively correlated with reasoning length, while reasoning length is strongly negatively correlated with pass rate. This mediation pattern can be summarized as higher difficulty $\rightarrow$ longer reasoning length $\rightarrow$ lower pass rate.
Table 7: Performance analysis by reasoning token length.
Token Total Passed Easy Medium Hard
0–5k 38 38 30/30 (100.0%) 8/8 (100.0%) 0/0 (–)
5k–10k 41 38 16/17 (94.1%) 14/16 (87.5%) 8/8 (100.0%)
10k–15k 41 32 10/11 (90.9%) 14/19 (73.7%) 8/11 (72.7%)
15k–20k 52 36 4/4 (100.0%) 16/16 (100.0%) 16/32 (50.0%)
20k–25k 36 15 1/1 (100.0%) 9/13 (69.2%) 5/22 (22.7%)
>>25k 60 10 0/0 (–) 2/14 (14.3%) 8/46 (17.4%)
Total 268 169 61/63 (96.8%) 63/86 (73.3%) 45/119 (37.8%)
Figure 7: Pass rate by token.

### v2

@@ 4 Experiment
Data and Training. We adopt GPT-o3-mini (OpenAI, 2025) for task formulation, DeepSeek-R1-0528 (DeepSeek-AI, 2025) and Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025) for solution sampling, and R1-0528 for test case generation. Statistics for SFT datasets are provided in Appendix C.2. For SFT, we set the learning rate at 5e-5, with a global batch size of 128 to train 8 epochs. For RL, the reward is defined as the fraction of passed tests among all given tests (detailed in Appendix A.2). The program executes in an isolated sandbox environment deployed with Redis, which supports optimized concurrent code testing (infrastructure details are provided in Appendix A.5). Training configurations and costs are supplemented in Appendix A.4.
@@
MiMo-SFT (Xiaomi et al., 2025) MiMo-Base 7B 500k Undisclosed avg@8 52.3 45.5
@@
X-Coder-Qwen2.5-SFT Qwen2.5-Coder-Instruct 7B 200k Syn avg@8 $\textbf{60.3}_{\pm 2.5}$ $\textbf{53.5}_{\pm 1.7}$
X-Coder-Qwen3-SFT Qwen3-8B-Base 8B 200k Syn avg@8 $\textbf{59.4}_{\pm 2.0}$ $\textbf{55.4}_{\pm 2.3}$
SFT-then-RL Baselines (Stage 2)
AceReason1.1 AceReason1.1-SFT 7B - Real avg@8 57.2 52.1
MiMo MiMo-SFT 7B 130k Undisclosed avg@8 57.8 49.3
Klear-Reasoner Klear-Reasoner-SFT 8B 106k Real avg@8 61.6 53.1
X-Coder-Qwen2.5 X-Coder-Qwen2.5-SFT 7B 40k Syn avg@8 $\textbf{62.9}_{\pm 1.8}$ $\textbf{55.8}_{\pm 1.9}$
X-Coder-Qwen3 X-Coder-Qwen3-SFT 8B 40k Syn avg@8 $\textbf{64.0}_{\pm 2.5}$ $\textbf{56.5}_{\pm 1.3}$
Figure 3: Left: Data scaling on LiveCodeBench v5. Right: Scaling unique tasks and scaling solutions per task.
Evaluation. We use LiveCodeBench v5 and v6 (Jain et al., 2024), the leading benchmarks for code reasoning. Using Qwen2.5-Coder-7B-Instruct and Qwen3-8B-Base as backbones, we report avg@8 scores (temperature 0.6, top-p 0.95, top-k 20) to ensure fair comparison with baselines. Beyond the primary setup, we assess the generalizability of synthetic training across the Llama model family and other code generation benchmarks in Appendix F.
4.1 Main Results
As shown in Table 1, during the SFT stage, X-Coder-SFT achieves an avg@8 pass rate of 60.3. X-Coder-SFT exhibits a clear advantage over 14B-based RL models (e.g., DeepCoder-Preview-14B, AReal-boba²-14B), despite those models being built on the stronger foundation R1-Distilled-Qwen. Relative to SFT-then-RL models, X-Coder further boosts its performance after RL, reaching 62.9. On Qwen3-Base, X-Coder attains an avg@8 pass rate of 64.0.
4.2 SFT Experiments and Analysis
Scaling Laws Hold for Synthetic Data. During the SFT stage, we investigate a central question: Can the SFT dataset be effectively scaled, and along which dimension should it be scaled more favorably? To explore this, we are inspired by AceReason-Nemotron 1.1 (Liu et al., 2025b) and expand the SFT dataset from two distinct perspectives: increasing the number of unique tasks and enlarging the number of solutions per task. We design six subsets (v1–v6): v1–v4 increase the number of unique tasks (32k, 64k, 128k, and 192k unique prompts, each with 1 solution), while v5–v6 expand the number of solutions per task (16k unique prompts with 4 solutions, and 8k unique prompts with 8 solutions). The results in Figure 3 reveal a promising scaling trend: performance improves monotonically as the dataset size increases (v1 $\to$ v4), rising steadily from 43.7% to 62.7%.
Table 2: Head-to-head Comparison with OpenCodeReasoning using Qwen2.5-Coder-7B-Instruct.
Model Avg. Pass@1 Pass@8 Easy Medium Hard
OpenCodeReasoning-Qwen 53.6 69.0 95.2 67.0 21.8
X-Coder-7B-SFT 60.3 77.0 96.8 73.3 37.8
@@ 4.2 SFT Experiments and Analysis
Synthetic Data Offers a Superior Alternative. We conduct a head-to-head comparison with OpenCodeReasoning (Ahmad et al., 2025), the largest reasoning dataset built upon real-world tasks. We train our dataset and OpenCodeReasoning using the same number of training tokens. The results are shown in Table 2. Our proposed dataset yields a 6.7-point improvement after SFT, with most gains coming from the medium and hard splits. These results demonstrate that fully synthetic data is not only a viable substitute for real-world data but also potentially offers a higher upper bound for scaling code reasoning. Moreover, training on fully synthetic data significantly reduces data contamination risks compared to models trained on real-world datasets (Appendix G).
@@ 4.3 RL Experiments and Analysis
Our investigation of the RL stage uncovers the following insights into its role and behavior:
RL as a Critical Performance Booster. RL fine-tuning is not merely a minor refinement but a critical optimization stage. As shown in Table 1, when applied to a converged SFT model using only code data, it yields a substantial 4.6% absolute gain in average pass-rate. This highlights RL’s unique capability to explore policy beyond the distribution of the initial supervised dataset.
Figure 4: Reward dynamics of weak and strong SFT models.
“Good-gets-Better” Principle. RL performance is tightly coupled to the strength of the SFT initializer. Using two SFT models trained on similar data distributions but with different LiveCodeBench scores as starting points, we observe in Figure 4 that, under identical RL settings, the stronger initializer consistently attains higher rewards.
A stronger SFT foundation enables the model to explore a more promising policy space and achieve a higher performance ceiling. This underscores the importance of a high-quality initial model as a prerequisite for effective RL.
RL’s Resilience to Noisy Supervision. Contrary to the common assumption that RL requires pristine reward signals, our experiments reveal a resilience to data imperfections during RL. The model also effectively benefits from synthetic test cases, suggesting that RL can be successfully deployed in scenarios with large-scale but imperfect feedback (Wang et al., 2020; Lv et al., 2025), significantly lowering the barrier to code RL data collection.
@@ 5 Ablation Study
Despite the strong performance of X-Coder, the determinants of high-quality synthetic data for SFT remain insufficiently understood. To elucidate these factors, we conduct a comprehensive ablation along four axes: (i) the effect of the proposed dual-verification strategy; (ii) the impact of distinct thinking types in solutions; (iii) a head-to-head comparison of tasks produced by the proposed synthesis versus those from open-source synthetic datasets; and (iv) data-selection strategies to identify patterns that shape downstream performance. Additionally, ablations on task styles and test generation strategies are provided in Appendix B.2.3 and Appendix D.2, respectively.
Table 3: Raw vs. verified solutions on LCB v5.
Dataset Size Verify LCB v5
Raw Solution 64k 47.0
Verified Solution 64k 53.4 (+6.4)
Dual-verification is Critical for Quality. We employ a dual-verification strategy to mitigate noise in stochastic sampling. Table 3 confirms its efficacy: training on verified solutions (64k tasks, Qwen2.5-Coder-7B-Instruct) significantly outperforms raw solutions. However, verifying 200k samples requires 1.6M CoT trajectories and 24M executions. This overhead establishes a clear trade-off. Although prior work (Li et al., 2025; Gandhi et al., 2025) indicates that models can learn effectively from unverified long-CoT data, making raw-solution training a more resource-efficient alternative, we contend that the significant cost of verification is justified to achieve the higher performance ceiling required for expert-level reasoning. Detailed validation effectiveness analysis is provided in Appendix E.3.
@@
Long Reasoning Chains are More Valuable but Converge Slower during Training. The length of CoT proves to be a critical factor for performance, with longer CoTs yielding superior results despite higher training costs. To demonstrate this, we compare the Qwen2.5-Coder-7B-Instruct trained on solutions generated by DeepSeek-R1-0528 (Long-CoT) and Qwen3-235B-A22B-Instruct-2507 (Short-CoT) for an identical set of tasks (200k).
@@
Table 5: Synthetic Data Comparison.
Method Evolution Adapted LCB v5
SelfCodeAlign-10k 27.1
Ours-10k 31.7
EpiCoder-64k 25.4
Ours-64k 46.3
Domain-Specific Evolution is Essential for Synthesis. As presented in Table 5, we compare the proposed strategy against two baselines to validate our methodological choices. First, compared to EpiCoder (using general-purpose prompts), the domain-adapted synthesis achieves a substantial 20.9% improvement. Since this method is built upon EpiCoder’s framework but injects competitive programming-oriented feature extraction and evolution, this gain directly validates the necessity of domain expertise for synthesizing high-complexity tasks. General-purpose synthesis is insufficient for expert-level domains. Second, compared to SelfCodeAlign (Wei et al., 2024) (where we applied the same domain adaptation), the proposed method still yields a 4.6% gain. This suggests that for competitive programming—a domain characterized by intricate logic and corner cases—the feature-based evolution strategy is more effective than concept composition at generating high-quality, diverse, and rigorous training data.
@@
Reasoning-Intensive Tasks are More Valuable. To investigate data utilization efficiency, we explore task selection strategies for competitive programming. Specifically, we evaluate three approaches: (1) difficulty-based selection, where GPT-4o-2411 assigns discrete difficulty scores to tasks, simulating the Codeforces rating system; (2) rationale-based selection, where DeepSeek-R1-0528 generates CoT reasoning for each task, and tasks that elicit longer reasoning traces are prioritized; and (3) random selection as a baseline. Each strategy independently samples a 50k-task subset from a 200k-task pool. Solutions are generated by Qwen3-235B-A22B-Instruct-2507, and models were trained for three epochs with a 16k context length.
As shown in Figure 5, tasks that induce longer CoT are regarded as more valuable training data for competitive programming, as they demand deeper reasoning and are potentially more challenging. This finding provides a practical heuristic for efficient data pruning in resource-constrained training scenarios.
Table 6: Distribution of failure cases for 16 rollouts on LiveCodeBench v5.
@@
Table 7: Performance analysis categorized by reasoning token length.
Token Range Total Passed Easy Medium Hard
0–10k $79$ $76$ $46/47\ (97.9\%)$ $22/24\ (91.7\%)$ $8/8\ (100.0\%)$
10k–20k $93$ $68$ $14/15\ (93.3\%)$ $30/35\ (85.7\%)$ $24/43\ (55.8\%)$
[…]
Total $\mathbf{268}$ $\mathbf{169}$ $\mathbf{61/63\ (96.8\%)}$ $\mathbf{63/86\ (73.3\%)}$ $\mathbf{45/119\ (37.8\%)}$
6 Discussion
Beyond metrics, understanding the mechanisms of failure and success is crucial for advancing code reasoning. In this section, we conduct a granular analysis of the model’s limitations and behavioral patterns. We examine the distribution of error types to pinpoint persistent bottlenecks, revealing an interplay between task difficulty, reasoning length, and success rates. Furthermore, we investigate the characteristics emerging from each training stage, covering the capability distillation in SFT as well as the test-time scaling behaviors and potential reward hacking risks in RL.
Reasoning Capability Remains the Primary Bottleneck. We classify failure cases into seven types: Wrong Answer (output mismatches the expected answer), Time Limit Exceeded, Memory Limit Exceeded, No Code Block Generated (truncated due to heavy reasoning before the final code is generated), Incomplete Code Block (partial code without closure), Function Signature Mismatch (incorrect function definition), and Syntax Error (complete code with syntax issues). The error distribution in Table 6 indicates that the primary bottleneck lies in reasoning capability, with most errors stemming from wrong answers. Two other major failure categories are No Code Block Generated and Time Limit Exceeded (TLE). Inspection of no-code samples reveals that all exceeded the 32k context window, resulting in truncated reasoning. The prevalent TLE errors further emphasize the necessity for Code LLMs to prioritize execution efficiency.